You are evaluating a capital project for equipment with a total installed cost of $750,000. The equipment has an estimated life of 30 years, with an expected salvage value at the end of the project of $50,000. The project will be depreciated via simplified straight-line depreciation method. In addition, a working capital investment of $5,000 is required. The project replaces an old piece of equipment which is currently in service and is fully depreciated, but has an expected after-tax salvage value of $12,000. After replacing the old equipment, cash savings from decreased operating expenses are expected to amount of $200,000 per year. The firm;s marginal tax rate is 40 percent and the project cost of capital is 10%. What is the net present value of this project? Round to the nearest penny. Do not include a dollar sign in your answer.
Answer #1

Statement showing NPV

Particulars 0 1-30 years 30 NPV
Cost of machine -750,000
WC required -5,000
After tax salvage value of old machine 12,000
Savings from reduction in cost 200,000
Depreciation 25,000
PBT 175,000
Tax @ 40% 70,000
PAT 105,000
Add: Depreciation 25,000
Annual cash flow 130,000
Release of WC 5000
After tax salvage value of new machine
(5000(1-tax rate))
=5000(1-0.4)
=50000(0.6) = 30,000
30000
Total cash flow -743,000 130,000 35000
PVIF @ 10% year 0 1
PVIF @ 10% year 30 0.0573
PVIF @ 10% year 30 9.4269
Present value ( Cash flow x PVIFA/PVIF) -743000 1225497 2006 484503

PVIF @10% 30 years = 1/(1+r)^n

=1/(1.1)^30 = 0.0573

PVIFA @ 10% 30 years = [1-(1/(1+r)^n)]/r

=[1-(1/(1.1)^30)]/0.1

=[1-0.0573]/0.1

=0.94269/0.1

=9.4269